Mathematics Examples

Simon Hood

February 22, 2006

Attempt to reproduce the equations in this document! You may need to look at the examples given in the links on the course page entitled Mathematics — More.

1 The math environment

Here are some straightforward examples: y = x^{2}, y = sin x.

2 The displaymath environment

Three examples:

\frac{x}{y} = z

(\frac{x}{y} + 3) = z

And now a harder one:

t = \frac{X_{1} − X_{2}}{\sqrt{P (\frac{1}{n_{1}} + \frac{1}{n_{2}} )}}.

3 The equation environment

Derivatives and greek letters:

ϵ\frac{d^{2}ψ}{dt^{2}} = −ψ.  (1)

4 The eqnarray environment

This environment is for use when you want to align equations with the equals sign or some other convenient point.

ϵ\frac{d^{2}ψ}{dt^{2}} = −ψ  (2)
q \frac{∂p}{∂t} = −∇.J_{p} − qR  (3)

5 Fine points

The first version is just plain wrong; the second is suitable for publication: log y = ∫_{0}^{∞} sin xdx

log y = ∫_{0}^{∞} sin x dx